OpenStudy (anonymous):

given that f(x)=-x+4 and g(x)=-2x-3, solve for f(g(x)) when x=2

OpenStudy (anonymous):

f(g(x)) = f(-2x-3) = -(-2x-3) + 4 = 2x+3+4 = 2x + 7 Keep in mind, that when we want f(-2x-3), we want the value of f, when x takes on a value of -2x-3, so just replace x with -2x-3.

OpenStudy (anonymous):

that's the answer ? @Easyaspi314

OpenStudy (anonymous):

yes, answer is 2x + 7. Any questions how I got it?

OpenStudy (anonymous):

no thanks !! (:

OpenStudy (anonymous):

cool.